OpenStudy (anonymous):

1/(2k-1) = (k+2)/25=4/(6k+2) Find K.!!! I could solve till this step, please help me proceed

OpenStudy (anonymous):

if there really 2 equal signs?

OpenStudy (anonymous):

Ya!! They have to be equated to find k! I am trying cross multiplication method but am not getting the answer!

OpenStudy (zzr0ck3r):

1/(2k-1) = (k+2)/25 (2k-1)(k+2)=25 2k^2+3k-27=0 note if the third equation is equal to the other two, then all we need to solve for is this^^^

OpenStudy (zzr0ck3r):

2(k^2+(3/2)k)= 27 2(k+(3/4))^2=27+18/16 (k+(3/4))^2= 27/2+18/32 k = +-sqrt(27/2+9/16)-(3/4)

OpenStudy (zzr0ck3r):

LHS evaluated at K 1/(2(sqrt(27/2+9/16)-(3/4))-1) = 1/5 RHS evaluated at K sqrt(27/2+9/16)-(3/4) = 1/5 I let K be the positive roots, you can check the negative, and since solving the 2nd,3rd equation is a quadratic, we know it will have no more than two solutions. so they are equal

OpenStudy (zzr0ck3r):

so all three are equal for k = +-sqrt(27/2+9/16)-(3/4)

OpenStudy (anonymous):

ThanQ!

OpenStudy (zzr0ck3r):

do you understand?

OpenStudy (zzr0ck3r):

I just completed the square to solve the equation, if you are not good at that, use what ever method you want.

OpenStudy (anonymous):

Yeah!! Now il take square root and solve for k! THanks!

OpenStudy (anonymous):

@zzr0ck3r - I am getting \[\sqrt{213/16}\] correct?

OpenStudy (anonymous):

@zzr0ck3r - Is that correct? I am not getting a round number!

OpenStudy (zzr0ck3r):

there are two numbers 3 and I forget the other, use a calculator, they are up there^^^

OpenStudy (anonymous):

Okay!

OpenStudy (zzr0ck3r):

sqrt(27/2+9/16)-(3/4)=3 -sqrt(27/2+9/16)-(3/4)=-4.5 sorry was on the phone.